Unchargeu. Widt three years? In eight years? In 12 years? In 13 years? Whats going on here? Illustrate your answers by gra

Hi,

Would you show me how to solve #19 problem with ONLY using a financial calculator?
Thank you very much for your time!
Answer #1

As bonds are priced at par, the ytm must be equal to coupon raate=6.5%

I will show for Bond Sam for 2% increase and 2% decrease

2% increase:
N=2*3=6
PMT=-6.5%*100/2=-3.25
I/Y=(6.5%+2%)/2=4.25%
FV=-100
CPT PV=94.800

% change=94.80/100-1=-5.200%

2% decrease:
N=2*3=6
PMT=-6.5%*100/2=-3.25
I/Y=(6.5%-2%)/2=2.25%
FV=-100
CPT PV=105.554

% change=105.554/100-1=5.554%